hello would you please help me with this problem? the tens digit of a two-digit number is 3 more then the units digit. the number is 8 more then six times the sum of the digits thank you for your help: hello would you please help me with this problem? the tens digit of a two-digit number is 3 more then the units digit. the number is 8 more then six times the sum of the digits thank you for your help
Answer by scott8148(2761)

"the tens digit of a two-digit number is 3 more then the units digit" __ t=u+3

"the number is 8 more then six times the sum of the digits" __ 10t+u=6(t+u)+8

distributing __ 10t+u=6t+6u+8 __ substituting __ 10(u+3)+u=6(u+3)+6u+8 __ distributing __ 10u+30+u=6u+18+6u+8

11u+30=12u+26 __ subtracting 11u+26 __ 4=u

substituting __ t=(4)+3 __ t=7

the number is 74